You are here: Foswiki>PartnerObserving Web>ItalyQueue2010>ObsInaf20100412 (14 Apr 2010, VincenzoTest)Edit Attach

INAF LBC+LUCIFER queue observing, 2010 April11-12

Observer: R. Speziali, K.Boutsia
Telescope Operator: D.Gonzales Huerta
Support Astronomer: D. Thompson
Instrument: LUCIFER1

Summary

Clear night with thin cirrus at the beginning. The seeing was ~1.0.
Lost 30min during MOS acquisition because Skycat fails to load new images.
At 09:40 (UT) rotator hits a limit. David G.H. had to go to the dome and fix it.
At 10:05 (UT) the dome is closed due to strong winds.

Data Summary

Project Objects Filter ExpTime Notes
Mannucci HDF-N 210zHJK + K 2hr arc+flat+telluric done
Mannucci Q1422 210zJHK + H 50min arc+flat+telluric done

Details

01:55 - Broad band Sky Flats
  • Ks flats 2secx10 (12000-9000) 0001-0010
  • Ks flats 2secx10 (9000-7000) 0011-0020
  • Ks flats 2secx10 (<5000) 0021-0030
  • Ks flats 2secx10 (<5000) 0031-0040
  • J flats 2secx10 (10000-6000) 0041-0050
  • J flats 2secx10 (<3000) 0051-0060

02:12 - Pointing Star High Elevation (WT10_303) 0063

02:18 - Pointing Star Low Elevation (ACT0097) 0064
  • -266.75 - 0.12*(1915[current x center]-1014[center of the chip])

02:22 - Pointing Star High Elevation (WT10_303) 0065

02:27 - Persson standard BS9145 for collimation (POSITION,ACTIVE)
  • guide star has also been used as a pointing star near the field to correct pointing 0066

02:36 - Collimation - preset to Mannucci HDFN
  • MOS acquisition 0067-0074
  • Skycat fails to load new images - loss of time!!!

03:30 - Mannucci HDFN K_1 SCI script 0075-0086

04:39 - Through slit image to see if sources are centered 0087-0088

04:46 - Mannucci HDFN K_2 SCI script 0089-0100

06:00 - Through slit 0101-0102
  • seeing 1.1
  • we stop HDFN K and we start calibration

06:05 - Calibration Unit in - Mannucci arc_HDF_K and Mannucci flat_HDF_K
  • Dark Ar 0103
  • ARC Ar 0104
  • Dark Xe 0105
  • ARC Xe 0106
  • Halo1 lamp off 0107-0109
  • Halo2 lamp on 0110-0112

06:15 - Slew to telluric for HDF
  • fail to find guide star
  • pointing correction - (WT10_385) 0113-0114

06:29 - HDF telluric acquisition 0115-0118

06:37 - HDF telluric integration script in K band 0119-0126

06:47 - Slew to Q1422 Mannucci
  • Telescope control crash - repeat MOS acquisition 0127-0132

07:33 - Mannucci Q1422 H_3 SCI script 0133-0141
  • script aborted after completing 9 out of 12 offsets
  • 10th offset aborted due to collimation error

08:38 - While resetting mirrors - move calibration unit in - Arc and Flat acquisition scripts
  • Dark Ar 0142
  • ARC Ar 0143
  • Dark Xe 0144
  • ARC Xe 0145
  • Halo1 lamp off 0146-0148
  • Halo2 lamp on 0149-0151

08:46 - Telluric star for Q1422 acquisition script
  • Abort - no guide star is found - correct pointing (ACT0279) 0152

08:50 - Slew to telluric for Q1422 acquisition 00153-0157

09:03 - Mannucci_tell_int_H for Q1422 0158-0165

09:12 - Mannucci_tell_int_K for Q1422 0168-0175 * verify that source is in the slit 0166-0167

09:18 - Slew back to Q1422 - MOS acquisition 0176-0178

09:33 - Q1422_H4 SCI script 0179-0180
  • no guiding during the second exposure

09:45 - Abort script - Lost preset
  • Rotator hit a limit
  • telescope operator had to go to the dome and fix it

10:05 - DOME CLOSED DUE TO WIND

12:50 - DoDarks.txt script - End of night

-- VincenzoTest - 12 Apr 2010
Topic revision: r6 - 14 Apr 2010, VincenzoTest
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback